# Categorical Flow Matching on Statistical Manifolds

### v1

@@ Abstract
We introduce Statistical Flow Matching (SFM), a novel and mathematically rigorous flow-matching framework on the manifold of parameterized probability measures inspired by the results from information geometry. We demonstrate the effectiveness of our method on the discrete generation problem by instantiating SFM on the manifold of categorical distributions whose geometric properties remain unexplored in previous discrete generative models. Utilizing the Fisher information metric, we equip the manifold with a Riemannian structure whose intrinsic geometries are effectively leveraged by following the shortest paths of geodesics. We develop an efficient training and sampling algorithm that overcomes numerical stability issues with a diffeomorphism between manifolds. Our distinctive geometric perspective of statistical manifolds allows us to apply optimal transport during training and interpret SFM as following the steepest direction of the natural gradient. Unlike previous models that rely on variational bounds for likelihood estimation, SFM enjoys the exact likelihood calculation for arbitrary probability measures. We manifest that SFM can learn more complex patterns on the statistical manifold where existing models often fail due to strong prior assumptions. Comprehensive experiments on real-world generative tasks ranging from image, text to biological domains further demonstrate that SFM achieves higher sampling quality and likelihood than other discrete diffusion or flow-based models.
@@ 1 Introduction
Recently, conditional flow matching (CFM) models [36] have achieved remarkable success in various generative domains including image generation [36, 18, 28], molecule [55, 54, 30] and protein design [62, 11], and sequence generation [56, 37, 13]. While attempts to generalize CFM and diffusion models to discrete categorical data have been made, they typically exert ad hoc assumptions on the structure of the discrete distribution. One group of work relies on stochastic jumps of Markov chains in either the discrete-time [6, 37, 1] or continuous-time setting [12, 51, 13] that discards the continuous nature of the underlying categorical distributions. Other work directly works with the probability simplex [7, 56] or the corresponding logit space [27, 38, 23] with potentially imperfect assumptions that fail to capitulate the underlying true geometry of the statistical manifold. Furthermore, likelihood is often approximated by variational bounds in previous discrete generative models due to the lack of tractable exact likelihood.
@@ 2.2 Conditional Flow Matching on Riemannian Manifold
Consider a smooth Riemannian manifold $\mathcal{M}$ with the Riemannian metric $g$, a probability path $p_{t}:[0,1]\to\mathcal{P(M)}$ is a curve of probability densities over $\mathcal{M}$. A flow $\psi_{t}:[0,1]\times\mathcal{M}\to\mathcal{M}$ is a time-dependent diffeomorphism defined by a time-dependent vector field $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ via the ordinary differential equation (ODE): $\frac{d}{dt}\psi_{t}(x)=u_{t}(\psi_{t}(x))$. The flow matching objective directly regresses the vector field $u_{t}$ with a time-dependent neural net $v(x_{t},t)$ where $x_{t}:=\psi_{t}(x)$. However, this objective is generally intractable. Both [36, 15] demonstrated that a tractable objective can be derived by conditioning on the delta measure $\delta_{x_{1}}$ at $t=1$ of the probability path. The Riemannian flow matching objective can be formulated as [15]
@@ 3.1 Statistical Manifold of Categorical Distributions
Consider the discrete sample space $\mathcal{X}=\{1,2,\dots,n\}$, an $n$-class categorical distribution over $\mathcal{X}$ can be parameterized by $n$ parameters $\mu_{1},\mu_{2},\dots,\mu_{n}$ such that $\sum_{i=1}^{n}\mu_{i}=1,\mu_{i}\geq 0$. In this way, the reference measure $\nu$ is the counting measure and the probability measure $\mu$ can be written as the convex combination of the canonical basis of delta measures $\{\delta^{i}\}_{i=1}^{n}$ over $\mathcal{X}$: $\mu=\sum_{i=1}^{n}\mu_{i}\delta^{i}$. Geometrically, this manifold can be visualized as the $(n-1)$-dimensional simplex $\Delta^{n-1}$. The geodesic distance between two categorical distributions is given in [48, 42] as
@@
The target space at a point $\mu$ can be identified with $T_{\mu}(\mathcal{P})=\{u|\sum_{i=1}^{n}u_{i}=0\}$ and the corresponding inner product at point $\mu$ is defined as
@@
where $\mathcal{P}^{\circ}$ denotes the interior of the statistical manifold $\mathcal{P}$. Note that the inner product is ill-defined on the boundary, causing numerical issues near the boundary. To circumvent this issue, we introduce the following diffeomorphism:
@@ 3.2 Statistical Flow Matching
We provide the analytical form of the exponential and logarithm maps on the statistical manifold of categorical distributions in Appendix A.3. Although it is possible to directly learn the vector field following the loss in Eq.2, such a direct parameterization has numerical issues near the boundary. As described in Sec.3.1, we apply the diffeomorphism $\pi$ in Eq.5 to derive a more stable training objective on the spherical manifold:
$\mathcal{L}_{\text{SFM}}=\mathbb{E}_{t\sim U[0,1],x_{0}\sim\pi(p_{0}(\mu)),x_{1}\sim\pi(q(\mu))}\left[\|v(x_{t},t)-u^{S}_{t}(x_{t}|x_{0},x_{1})\|_{2}^{2}\right]$ (8)
where $p_{0}$ is the prior noise distribution over $\mathcal{P}$ and $q$ is the data distribution. $v:S^{n-1}_{+}\times[0,1]\to TS^{n-1}_{+}$ is a learnable time-dependent vector field network that maps the interpolant $x_{t}$ on the unit sphere to a tangent vector in the tangent bundle $TS^{n-1}_{+}$. The ground truth conditional vector field $u^{S}$ is calculated using the exponential and logarithms maps on the sphere (Appendix A.2). The overall training and inference scheme is visualized in Fig.2 and described in Appendix Alg. 2 - 3.
@@
Figure 2: Statistical flow matching (SFM) framework. (a) During training (Sec.3.2), probability measures on $\mathcal{P}$ are mapped to $\smash{S^{n-1}_{+}}$ via diffeomorphism $\pi$ to compute the time-dependent vector field (in red). During inference, the learned vector field generates the trajectory on $\smash{S^{n-1}_{+}}$ and we map the outcome of ODE back to $\mathcal{P}$ (in blue). (b) In the NLL calculation for one-hot examples (Sec.3.5), the probability density is marginalized over a small neighborhood of some delta measure to avoid undefined behaviors at the boundary (in green).
@@ 3.3 Optimization View of Statistical Flow Matching
We further provide an interpretation of our proposed statistical flow matching framework from the perspective of an optimization problem. For a generative model on a statistical manifold, we want to minimize the discrepancy between the target distributions and the generated distributions. Naturally, the Kullback-Leibler divergence $D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))$ can be used as a measure of statistical distance. We note that the Fisher information metric can be obtained as the Hessian of the KL divergence with respect to the parameter $\theta$. This can be demonstrated by the fact that the KL divergence reaches the global minimum of 0 at $\theta=\theta_{1}$, so all first-order partial derivatives are zero and the Hessian is positive semi-definite. Therefore, Taylor expansion of KL divergence at $\theta_{1}$ gives
$D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))\approx\frac{1}{2}\sum_{jk}\Delta\theta_{j}\Delta\theta_{k}\left.\frac{\partial^{2}}{\partial\theta_{j}\partial\theta_{k}}D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))\right|_{\theta=\theta_{1}}=\frac{1}{2}\sum_{jk}\Delta\theta_{j}\Delta\theta_{k}g_{jk}(\theta_{1})$ (10)
Therefore, the Fisher information metric defines a second-order optimization scheme for the KL divergence by following the “steepest” direction of the Hessian. In fact, this idea has already been explored by previous work on optimization known as natural gradient descent [2, 3] and has been applied to training machine learning models [46, 40, 44]. Instead of optimizing along the normal gradient, the natural gradient is defined as ~$\tilde{\nabla}\mathcal{L}=F^{-1}\nabla\mathcal{L}$ where $F$ is the Fisher information matrix estimated from a batch of sampled data. Results established in [2, 40] demonstrated that stochastic natural gradient descent is asymptotically “Fisher efficient” and is indeed the steepest descent direction in the manifold of distributions where distance is measured in small local neighborhoods by the KL divergence. Following the geodesic defined by the Fisher information metric, our SFM framework also shares these benefits with an additional advantage of analytical expressions for geodesics, as we focus on the family of categorical distributions instead of general statistical models.
@@ 3.4 Optimal Transport on Statistical Manifold
The geometric viewpoint of the statistical manifold offers a continuous and differentiable formulation of generative modeling and also provides a robust way to measure the distance between two categorical distributions via the well-defined geodesic distance in Eq.(3). In contrast, the Markov chain-based methods cannot establish a robust distance measure due to the stochastic jumps between discrete states. Inspired by the optimal transport formulation in previous work [43, 20, 62], we naturally extend it to our statistical setting in which the cost is defined by averaging the statistical distances over data dimensions as $d_{\text{cat}}(X,Y)=\frac{1}{D}\sum_{k=1}^{D}d_{\text{cat}}(\mu^{(k)},\nu^{(k)})$ for $X=\{\mu^{(k)}\}_{k=1}^{D},Y=\{\nu^{(k)}\}_{k=1}^{D}$. An optimal assignment of the initial noises to the target data distributions can potentially lead to more efficient training, which we demonstrated empirically in our ablation studies.
@@ 3.5 Exact Likelihood Calculation
where $\mathrm{div}_{g}$ is the Riemannian divergence and $v_{t}(x_{t}):=v(x_{t},t)$ is the time-dependent vector field. In this way, the push-forward probability measures $p_{t}(x_{t})$ defined via the learned flow can be obtained as the integral of the Riemannian divergence back through time on the simulated trajectory $x_{t}$. Following [10, 7], we define the ODE log-likelihood as:
@@
where the trajectory $x_{t}$ is obtained by solving the differential equation $\frac{\partial}{\partial t}x_{t}=v_{t}(x_{t})$ reverse through time with the initial condition $x_{1}$ at $t=1$. In this way, we have $\log p(x_{1})=\log p^{\text{ODE}}+\log p_{0}(x_{0})$ where $\log p_{0}(x_{0})$ is the log-likehood of the prior distribution at data point $x_{0}$. In practice, we follow previous work [7] to use Hutchinson’s trace estimator [29] to efficiently obtain an unbiased estimation of the divergence using standard normal random variables. To further account for the transform $\pi$ from $\mathcal{P}$ to $S^{n-1}_{+}$ and the reverse transform $\pi^{-1}$, additional change of measure formulae need to be applied. Consider the push-forward of the probability measure $p$ over $\mathcal{P}$ under the diffeomorphism $\pi$ defined by $\pi_{*}p(V):=p(\pi^{-1}V)$, we have the change of measure identity $\pi_{*}p(\pi(\mu))|\det d\pi(\mu)|=p(x)$ for $x=\pi(\mu)$. Therefore, adding the two log-determinant of the push-forward measures, the log-likelihood can be formulated as
@@
The above formula is well-defined for all interior points of $\mathcal{P}$, but the change of measure terms are undefined on the boundary. This can be understood as the fact that discrete likelihoods can be arbitrarily high [58]. Following [7], we derive a variational lower bound for the likelihood as the marginalized probability over the small neighborhood of the delta measure $\mu_{1}=\delta$ at the boundary:
@@ 4 Experiments
Our proposed SFM framework can be leveraged to approximate arbitrary distribution over $\mathcal{P}$, i.e., any distribution over the parameterized family of probability measures. We first demonstrate this with the toy example of a Swiss roll distribution on the probability simplex. We then conduct extensive experiments on real-world datasets for discrete generation across various domains including computer vision, natural language processing, and bioinformatics to demonstrate the effectiveness of our proposed model. We train and evaluate two different versions of the SFM framework: SFM w/ OT for our proposed model with optimal transport during training and SFM w/o OT for the model without optimal transport. We also implement a linear flow matching model on the probability simplex using the loss in Eq.(9) as an additional baseline, for which we dub LinearFM. For discrete data, we always use Eq.(14) to obtain finite negative log-likelihood (NLL). More details regarding model architectures and the experimental setup can be found in Appendix C.

### v2

@@ Abstract
We introduce Statistical Flow Matching (SFM), a novel and mathematically rigorous flow-matching framework on the manifold of parameterized probability measures inspired by the results from information geometry. We demonstrate the effectiveness of our method on the discrete generation problem by instantiating SFM on the manifold of categorical distributions whose geometric properties remain unexplored in previous discrete generative models. Utilizing the Fisher information metric, we equip the manifold with a Riemannian structure whose intrinsic geometries are effectively leveraged by following the shortest paths of geodesics. We develop an efficient training and sampling algorithm that overcomes numerical stability issues with a diffeomorphism between manifolds. Our distinctive geometric perspective of statistical manifolds allows us to apply optimal transport during training and interpret SFM as following the steepest direction of the natural gradient. Unlike previous models that rely on variational bounds for likelihood estimation, SFM enjoys the exact likelihood calculation for arbitrary probability measures. We manifest that SFM can learn more complex patterns on the statistical manifold where existing models often fail due to strong prior assumptions. Comprehensive experiments on real-world generative tasks ranging from image, text to biological domains further demonstrate that SFM achieves higher sampling quality and likelihood than other discrete diffusion or flow-based models. Our code is available at https://github.com/ccr-cheng/statistical-flow-matching.
11footnotetext: Equal contribution.
@@ 1 Introduction
Recently, conditional flow matching (CFM) models [37] have achieved remarkable success in various generative domains including image generation [37, 17, 28], molecule [58, 57, 30] and protein design [67, 10, 35], and sequence generation [59, 38, 12]. While attempts to generalize CFM and diffusion models to discrete categorical data have been made, they typically exert ad hoc assumptions on the structure of the discrete distribution. One group of work relies on stochastic jumps of Markov chains in either the discrete-time [6, 38, 1] or continuous-time setting [11, 54, 12] that discards the continuous nature of the underlying categorical distributions. Other work directly works with the probability simplex [7, 59] or the corresponding logit space [27, 39, 23] with potentially imperfect assumptions that fail to capture the underlying true geometry of the statistical manifold. Furthermore, likelihood is often approximated by variational bounds in previous discrete generative models due to the lack of tractable exact likelihood.
@@ 2.2 Conditional Flow Matching on Riemannian Manifold
Consider a smooth Riemannian manifold $\mathcal{M}$ with the Riemannian metric $g$, a probability path $p_{t}:[0,1]\to\mathcal{P(M)}$ is a curve of probability densities over $\mathcal{M}$. A flow $\psi_{t}:[0,1]\times\mathcal{M}\to\mathcal{M}$ is a time-dependent diffeomorphism defined by a time-dependent vector field $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ via the ordinary differential equation (ODE): $\frac{d}{dt}\psi_{t}(x)=u_{t}(\psi_{t}(x))$. The flow matching objective directly regresses the vector field $u_{t}$ with a time-dependent neural net $v(x_{t},t)$ where $x_{t}:=\psi_{t}(x)$. However, this objective is generally intractable. Both [37, 14] demonstrated that a tractable objective can be derived by conditioning on the target data $x_{1}$ at $t=1$ of the probability path. The Riemannian flow matching objective can be formulated as [14]
@@ 3.1 Statistical Manifold of Categorical Distributions
Consider the discrete sample space $\mathcal{X}=\{1,2,\dots,n\}$, an $n$-class categorical distribution over $\mathcal{X}$ can be parameterized by $n$ parameters $\mu_{1},\mu_{2},\dots,\mu_{n}$ such that $\sum_{i=1}^{n}\mu_{i}=1,\mu_{i}\geq 0$. In this way, the reference measure $\nu$ is the counting measure and the probability measure $\mu$ can be written as the convex combination of the canonical basis of Dirac measures $\{\delta^{i}\}_{i=1}^{n}$ over $\mathcal{X}$: $\mu=\sum_{i=1}^{n}\mu_{i}\delta^{i}$. Geometrically, this manifold can be visualized as the $(n-1)$-dimensional simplex $\Delta^{n-1}$. The geodesic distance between two categorical distributions is given in [50, 43] as
@@
The tangent space at a point $\mu$ can be identified with $T_{\mu}(\mathcal{P})=\{u|\sum_{i=1}^{n}u_{i}=0\}$ and the corresponding inner product at point $\mu$ is defined as
@@
where $\mathcal{P}^{\circ}$ denotes the interior of the statistical manifold $\mathcal{P}$. Note that the inner product is ill-defined on the boundary, causing numerical issues near the boundary. To circumvent this issue, we introduce the following diffeomorphism
@@ 3.2 Statistical Flow Matching
We provide the analytical form of the exponential and logarithm maps on the statistical manifold of categorical distributions in Appendix A.3. Although it is possible to directly learn the vector field following the loss in Eq.(2), such a direct parameterization has numerical issues near the boundary. As described in Sec.3.1, we apply the diffeomorphism $\pi$ in Eq.(5) to derive a more stable training objective on the spherical manifold:
$\mathcal{L}_{\text{SFM}}=\mathbb{E}_{t\sim U[0,1],x_{0}\sim\pi_{*}(p_{0}(\mu)),x_{1}\sim\pi_{*}(q(\mu))}\left[\|v(x_{t},t)-u^{S}_{t}(x_{t}|x_{0},x_{1})\|_{2}^{2}\right]$ (8)
where $p_{0}$ is the prior noise distribution over $\mathcal{P}$ and $q$ is the data distribution; $\pi_{*}$ denotes the standard pushforward measure. $v:S^{n-1}_{+}\times[0,1]\to TS^{n-1}_{+}$ is a learnable time-dependent vector field network that maps the interpolant $x_{t}$ on the unit sphere to a tangent vector in the tangent bundle $TS^{n-1}_{+}$. The ground truth conditional vector field $u^{S}$ is calculated using the exponential and logarithms maps on the sphere (Appendix A.2). The overall training and inference scheme is visualized in Fig.2 and described in Alg.2 and 3 in Appendix C.
@@
Figure 2: Statistical flow matching (SFM) framework. (a) During training (Sec.3.2), probability measures on $\mathcal{P}$ are mapped to $\smash{S^{n-1}_{+}}$ via diffeomorphism $\pi$ to compute the time-dependent vector field (in red). During inference, the learned vector field generates the trajectory on $\smash{S^{n-1}_{+}}$ and we map the outcome of ODE back to $\mathcal{P}$ (in blue). (b) In the NLL calculation for one-hot examples (Sec.3.5), the probability density is marginalized over a small neighborhood of some Dirac measure to avoid undefined behaviors at the boundary (in green).
@@ 3.3 Optimization View of Statistical Flow Matching
We further provide an interpretation of our proposed statistical flow matching framework from the perspective of an optimization problem. From the optimization viewpoint, for a generative model on a statistical manifold, we want to minimize the discrepancy between the target distributions and the generated distributions. Naturally, the Kullback-Leibler divergence $D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))$ can be used as a measure of statistical distance. We note that the Fisher information metric can be obtained as the Hessian of the KL divergence with respect to the parameter $\theta$. This can be demonstrated by the fact that the KL divergence reaches the global minimum of 0 at $\theta=\theta_{1}$, so all first-order partial derivatives are zero and the Hessian is positive semi-definite. Therefore, Taylor expansion of KL divergence at $\theta_{1}$ with $\Delta\theta=\theta-\theta_{1}$ gives
$D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))$ $\approx\frac{1}{2}\sum_{jk}\Delta\theta_{j}\Delta\theta_{k}\left.\frac{\partial^{2}}{\partial\theta_{j}\partial\theta_{k}}D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))\right|_{\theta=\theta_{1}}$ (10)
$=\frac{1}{2}\sum_{jk}\Delta\theta_{j}\Delta\theta_{k}g_{jk}(\theta_{1})=\frac{1}{2}\|\Delta\theta\|_{g}^{2}$
From the geometric viewpoint, the geodesic, by definition, is a (locally) length-minimizing curve with respect to the corresponding Riemannian metric. Therefore, by following the direction of the vector field that decreases the geodesic element $ds^{2}=\|d\theta\|^{2}_{g}\approx\|\Delta\theta\|^{2}_{g}$, we are indeed following the steepest direction that minimizes the local KL divergence. In this sense, the Fisher information metric defines a second-order optimization scheme for the KL divergence by following the “steepest” direction of the Hessian. Indeed, The steepest direction $\Delta\theta$ that decreases the KL divergence is known as the natural gradient in the existing literature [2, 3] and has been explored in optimization known as natural gradient descent[47, 41, 45]. Instead of optimizing along the normal gradient, the natural gradient is defined as ~$\tilde{\nabla}\mathcal{L}=F^{-1}\nabla\mathcal{L}$ where $F$ is the Fisher information matrix estimated from a batch of sampled data. Results established in [2, 41] demonstrated that stochastic natural gradient descent is asymptotically “Fisher efficient” and is indeed the steepest descent direction in the manifold of distributions where distance is measured in small local neighborhoods by the KL divergence. Following the geodesic defined by the Fisher information metric, our SFM framework also shares these benefits with an additional advantage of analytical expressions for geodesics, as we focus on the family of categorical distributions instead of general statistical models. Such a theoretical connection may contribute to the better performance of SFM.
@@ 3.4 Optimal Transport on Statistical Manifold
The geometric viewpoint of the statistical manifold offers a continuous and differentiable formulation of generative modeling and also provides a robust way to measure the distance between two categorical distributions via the well-defined geodesic distance in Eq.(3). In contrast, the Markov chain-based methods cannot establish a robust distance measure due to the stochastic jumps between discrete states. Inspired by the optimal transport formulation in previous work [44, 20, 67, 62], we naturally extend it to our statistical setting in which the cost is defined by averaging the statistical distances over data dimensions as $d_{\text{cat}}(X,Y)=\frac{1}{D}\sum_{k=1}^{D}d_{\text{cat}}(\mu^{(k)},\nu^{(k)})$ for $X=\{\mu^{(k)}\}_{k=1}^{D},Y=\{\nu^{(k)}\}_{k=1}^{D}$. An optimal assignment of the initial noises to the target data distributions can potentially lead to more efficient training, which we demonstrated empirically in our ablation studies.
@@ 3.5 Exact Likelihood Calculation
where $\mathrm{div}_{g}$ is the Riemannian divergence and $v_{t}(x_{t}):=v(x_{t},t)$ is the time-dependent vector field. In this way, the pushforward probability measures $p_{t}(x_{t})$ defined via the learned flow can be obtained as the integral of the Riemannian divergence back through time on the simulated trajectory $x_{t}$. Following [9, 7], we define the ODE log-likelihood as the change of the log-likelihood as:
@@
where the trajectory $x_{t}$ is obtained by solving the differential equation $\frac{\partial}{\partial t}x_{t}=v_{t}(x_{t})$ reverse through time with the initial condition $x_{1}$ at $t=1$. In this way, we have $\log p(x_{1})=\log p^{\text{ODE}}+\log p_{0}(x_{0})$ where $\log p_{0}(x_{0})$ is the log-likehood of the prior distribution at data point $x_{0}$. In practice, we follow previous work [7] to use Hutchinson’s trace estimator [29] to efficiently obtain an unbiased estimation of the divergence using standard normal random variables. To further account for the transform $\pi$ from $\mathcal{P}$ to $S^{n-1}_{+}$ and the reverse transform $\pi^{-1}$, additional change of measure formulae need to be applied. Consider the pushforward of the probability measure $P$ over $\mathcal{P}$ under the diffeomorphism $\pi$ defined by $\pi_{*}P(V):=P(\pi^{-1}V)$, we have the change of measure identity $\pi_{*}P(\pi(\mu))|\det d\pi(\mu)|=P(x)$ for $x=\pi(\mu)$. Therefore, by adding the two log-determinant of the pushforward measures, the log-likelihood can be formulated as
@@
The above formula is well-defined for all interior points of $\mathcal{P}$, but the change of measure terms are undefined on the boundary. This can be understood as the fact that discrete likelihoods can be arbitrarily high [61]. Following [7], we derive a variational lower bound for the likelihood as the marginalized probability over the small neighborhood of a Dirac measure $\mu_{1}=\delta$ at the boundary:
@@ 4 Experiments
Our proposed SFM framework can be leveraged to approximate arbitrary distribution over $\mathcal{P}$, i.e., any distribution over the parameterized family of probability measures. We first demonstrate this with the toy example of a Swiss roll distribution on the probability simplex. We then conduct extensive experiments on real-world datasets for discrete generation across various domains including computer vision, natural language processing, and bioinformatics to demonstrate the effectiveness of our proposed model. We train and evaluate two different versions of the SFM framework: SFM w/ OT for our proposed model with optimal transport during training and SFM w/o OT for the model without optimal transport. A naive approach that directly works with the exponential and logarithm maps on the statistical manifold without the diffeomorphism (SFM w/o $\boldsymbol{\pi}$) is also compared in the toy example. We also implement a linear flow matching model on the probability simplex using the loss in Eq.(9) as an additional baseline, for which we dub LinearFM. For discrete data, we always use Eq.(14) to obtain finite negative log-likelihood (NLL). More details regarding model architectures and the experimental setup can be found in Appendix C.